OpenStudy (anonymous):

-1 4/9 - (-2 5/18) perform the indicated operation.

OpenStudy (anonymous):

(-9/9)+(4/9)+(36/18)-(5/18) like this? if so, then (-18/18)+(8/18)+(36/18)-(5/18) = 21/18 = 7/6 7/6 being the simplest form.

OpenStudy (anonymous):

i got 15/8...reduced to 5/6? check work.

OpenStudy (anonymous):

Rationalize the mixed numbers -(1+ 4/9) and -(2+5/18). 9 times 1 + 4 over 9. 18 times 2 plus 5 over 18, respectively. \[ \left(-\frac{13}{9}\right)-\left(-\frac{41}{18}\right)=\left(-\frac{13}{9}\right)+\left(\frac{41}{18}\right)=\left(-\frac{26}{18}\right)+\left(\frac{41}{18}\right)=\frac{15}{18}=\frac{5}{6} \]